# Board Game Arena: A Framework and Benchmark for Assessing Large Language Models via Strategic Play

### v1

@@
Board Game Arena: A Framework and Benchmark for Assessing Large Language Models via Strategic Play
@@
(August 5, 2025)
The Board Game Arena library provides a framework for evaluating the decision making abilities of large language models (LLMs) through strategic board games implemented in Google’s OpenSpiel library. The framework enables systematic comparisons between LLM-based agents and other agents (random, human, reinforcement learning agents, etc.) in various game scenarios by wrapping multiple board and matrix games and supporting different agent types. It integrates API access to models via LiteLLM, local model deployment via vLLM, and offers distributed execution through Ray. This paper summarizes the structure, key characteristics, and motivation of the repository, highlighting how it contributes to the empirical evaluation of the reasoning of LLM and game-theoretic behavior. The complete library documentation can be found at https://board-game-arena.readthedocs.io/.
@@ 1 Introduction
Recent advances in large language models have spurred interest in evaluating their reasoning and planning abilities beyond standard natural language benchmarks. Strategic games offer a controlled environment where agents must plan, adapt and anticipate opponents’ moves, making them a natural testbed for decision making. Board Game Arena addresses this need by providing a unified framework for playing and analysing games such as Tic–Tac–Toe, Connect Four, Kuhn Poker and matrix games like Prisoner’s Dilemma. It uses Google‑DeepMind’s OpenSpiel library as the underlying game engine. OpenSpiel is an open‑source collection of environments and algorithms for reinforcement learning and search/planning in games; it supports single‑ and multi‑agent, zero‑sum and general‑sum games with perfect and imperfect information and includes tools to analyse learning dynamics. By building on this engine, the framework exposes a flexible API to configure games, agents and evaluation settings.
@@ 5.3 Results
To analyse the reasoning processes of LLMs in board game environments, we propose a categorization framework that labels model-generated justifications according to distinct reasoning patterns as shown in Figure 1. These include positional reasoning, which emphasizes spatial control of the board (e.g., references to the ”center column” or ”corner”); opponent modeling, where the model anticipates or reacts to the opponent’s strategy; and blocking, which denotes defensive actions aimed at preventing the opponent’s success. Additional categories include winning logic, which refers to direct strategies for achieving victory (e.g., ”winning move”, ”fork”); heuristic-based reasoning, characterized by general evaluative language (e.g., ”best move”, ”advantageous”); and rule-based reasoning, which appeals to known strategies or game rules. Finally, we include a random/unjustified category for responses lacking clear strategic motivation. By associating these types of reasoning with specific lexical cues, this classification enables a structured analysis of LLM decision making, highlighting patterns, strengths, and potential shortcomings in model reasoning across varied game scenarios.
Figure 1: Categorization scheme for analyzing LLM reasoning in board games. The table defines reasoning types based on patterns in the model’s output, with associated keywords used for identification and classification. This categorization enables structured analysis of model-generated justifications during gameplay.
Once we have mapped groups of words to a reasoning type, we analyse the reasoning traces provided by each LLM at game play. We use a word matching pattern to extract the key words off each LLM’s reasoning trace. This allows us to cluster reasoning traces by their type. As an example, we provide the analysis of the reasoning traces of the llm_codegemma_7b_it model across three different games—Connect Four, Kuhn Poker, and Tic-Tac-Toe—by categorizing its justifications according to the predefined reasoning taxonomy described in table 1. The pie charts reveal distinct patterns of reasoning that vary with game structure. In Connect Four 2, the model predominantly relied on blocking reasoning (75%), indicating a strong focus on defensive play to prevent the opponent from winning, while positional reasoning (25%) was used to justify spatial control. In Kuhn Poker 3, the reasoning was exclusively framed in terms of winning logic (100%), reflecting the strategic and payoff-driven nature of the game where positioning and blocking are irrelevant. In contrast, Tic-Tac-Toe 4showed a combination of blocking (70%) and heuristic-based reasoning (30%), suggesting that the model combined reactive defense with general evaluative principles such as optimality of movement. These findings indicate that the model adapts its reasoning style to game dynamics, favoring defensive reasoning in spatial games and goal-oriented logic in strategic card games.
Figure 2: Distribution of reasoning types used by llm_codegemma_7b_it in Connect Four. The model predominantly relied on blocking (75%), with some use of positional reasoning (25%), indicating a focus on defensive moves and spatial control.
Figure 3: Reasoning type distribution for llm_codegemma_7b_it in Kuhn Poker. All reasoning traces (100%) were classified as winning logic, reflecting the game’s strategic and payoff-oriented nature.
Figure 4: Reasoning type distribution for llm_codegemma_7b_it in Tic-Tac-Toe. The model relied primarily on blocking (70%) and to a lesser extent on heuristic-based reasoning (30%), suggesting a mix of defensive tactics and general strategy.
@@ 7 Related Research
Collectively, these benchmarks indicate that while LLMs excel at many language tasks, strategic gameplay remains challenging. They highlight the importance of dynamic, multi‑agent and interactive environments for evaluating reasoning, planning and social skills, and suggest motivation for frameworks such as OpenSpiel Lanctot et al. (2019) or Board Game Arena.

### v2

# Game Reasoning Arena: A Framework and Benchmark for Assessing Reasoning Capabilites of Large Language Models via Game Play

@@
Game Reasoning Arena: A Framework and Benchmark for Assessing Reasoning Capabilites of Large Language Models via Game Play
@@
(August 10, 2025)
The Game Reasoning Arena library provides a framework for evaluating the decision making abilities of large language models (LLMs) through strategic board games implemented in Google’s OpenSpiel library. The framework enables systematic comparisons between LLM‑based agents and other agents (random, heuristic, reinforcement learning agents, etc.) in various game scenarios by wrapping multiple board and matrix games and supporting different agent types. It integrates API access to models via liteLLM, local model deployment via vLLM, and offers distributed execution through Ray. This paper summarises the library’s structure, key characteristics, and motivation of the repository, highlighting how it contributes to the empirical evaluation of the reasoning of LLM and game‑theoretic behaviour. The complete library documentation can be found at https://game-reasoning-arena.readthedocs.io/ and the library is hosted at https://github.com/SLAMPAI/game_reasoning_arena.
@@ 1 Introduction
Recent advances in large language models have spurred interest in evaluating their reasoning and planning abilities beyond standard natural language benchmarks. Strategic games offer a controlled environment where agents must plan, adapt and anticipate opponents’ moves, making them a natural testbed for decision maGame Reasoning Arenaaddresses this need by providing a unified framework for playing and analysing games such as Tic–Tac–Toe, Connect Four, Kuhn Poker and matrix games like Prisoner’s Dilemma. It uses Google‑DeepMind’s OpenSpiel library as the underlying game engine. OpenSpiel is an open‑source collection of environments and algorithms for reinforcement learning and search/planning in games; it supports single‑ and multi‑agent, zero‑sum and general‑sum games with perfect and imperfect information and includes tools to analyse learning dynamics. By building on this engine, the framework exposes a flexible API to configure games, agents and evaluation settings.
@@ 5.3 Results
To analyse the reasoning processes of LLMs in board game environments, we propose a categorization framework that labels model-generated justifications according to distinct reasoning patterns as shown in Table 1. These include positional reasoning, which emphasizes spatial control of the board (e.g., references to the ”center column” or ”corner”); opponent modeling, where the model anticipates or reacts to the opponent’s strategy; and blocking, which denotes defensive actions aimed at preventing the opponent’s success. Additional categories include winning logic, which refers to direct strategies for achieving victory (e.g., ”winning move”, ”fork”); heuristic-based reasoning, characterized by general evaluative language (e.g., ”best move”, ”advantageous”); and rule-based reasoning, which appeals to known strategies or game rules. Finally, we include a random/unjustified category for responses lacking clear strategic motivation. By associating these types of reasoning with specific lexical cues, this classification enables a structured analysis of LLM decision making, highlighting patterns, strengths, and potential shortcomings in model reasoning across varied game scenarios.
We apply this classification framework to large-scale gameplay logs from multiple LLMs and games, producing the following multi-perspective analysis.
Table 1: Reasoning types with descriptions and representative words.
Reasoning Type Description Words
Positional Focuses on spatial board control (e.g. “center column”) center column, “center square”, “corner”, “edge”
Opponent modeling Refers to or anticipates opponent strategy “opponent”, “they are trying”, “their strategy”, “their move”
Blocking Refers to countering opponent threats “block”, “prevent”, “stop opponent”, “avoid opponent”, “counter”
Winning logic Direct mentions of creating win conditions “win”, “winning move”, “connect”, “fork”, “threat”, “chance of winning”
Heuristic-based General heuristics like “center is best” “best move”, “most likely”, “advantageous”, “better chance”
Rule-based Based on known strategies (forks, chains) according to, “rule”, “strategy”
Random/unjustified Reasoning lacks strategic justification “random”, “guess”
5.3.1 Temporal Adaptation Across Games (Radar Overview)
Figure 1 summarises temporal reasoning patterns for llm_litellm_groq_llama3_70b_8192 across seven games. Each polygon corresponds to a game; axes are reasoning types and radii denote normalised frequencies over the course of play. We observe opponent-focused behaviour in tic_tac_toe and prisoners_dilemma, positional emphasis in connect_four, rule-based regularity in matching_pennies, and a larger uncategorised share in matrix_pd.
Figure 1: Temporal reasoning-type patterns across games for llm_litellm_groq_llama3_70b_8192.
5.3.2 Aggregate Per-Game Profiles (Stacked Bars and By-Game Views)
Figures 26 report the overall per-game distribution of reasoning types. Stacked bars offer a compact percentage view for each game; the by-game plots show the same information with separate bars per category.
The 70B model exhibits richer variety and stronger opponent modelling in tic_tac_toe and matching_pennies, positional play in connect_four, and uncategorised traces in the matrix games. The 8B model shows heavier reliance on few categories and more frequent uncategorised segments.
Figure 2: Stacked reasoning percentages by game: llm_litellm_groq_llama3_70b_8192.
Figure 3: Stacked reasoning percentages by game: llm_litellm_groq_llama3_8b_8192.
Figure 4: By-game reasoning percentages: Meta LLaMA 3.1 8B Instruct Turbo.
Figure 5: By-game reasoning percentages: llm_litellm_groq_llama3_70b_8192.
Figure 6: By-game reasoning percentages: llm_litellm_groq_llama3_8b_8192.
5.3.3 Within-Game Evolution (Same Game, Different Turns)
We bin turns and compute per-bin reasoning proportions to detect shifts in strategy within a single game. Heatmaps emphasise concentration and switching; stacked bars show the mixture per bin. In tic_tac_toe, both models shift towards pure Blocking in the endgame, but their openings differ—one starts with Winning Logic then Opponent Modeling, while the other begins with strong Opponent Modeling and a positional midgame.
Figure 7: Within-game reasoning heatmap (tic_tac_toe): llm_litellm_groq_llama_3_1_8b_instant.
Figure 8: Turn-binned proportions (tic_tac_toe): llm_litellm_groq_llama_3_1_8b_instant.
Figure 9: Within-game reasoning heatmap (tic_tac_toe): llm_litellm_groq_llama3_8b_8192.
Figure 10: Turn-binned proportions (tic_tac_toe): llm_litellm_groq_llama3_8b_8192.
5.3.4 Per-Game Pie Charts (Detailed Breakdown Examples)
We include game-specific pies to illustrate the detailed mix behind the aggregates. In tic_tac_toe the instant 8B model balances Blocking (50%), Winning Logic (25%), and Opponent Modeling (25%). The 8B 8192 variant diversifies further, and in kuhn_poker the 70B model emphasises Opponent Modeling with smaller rule-based and uncategorised shares.
Figure 11: Pie breakdown (tic_tac_toe): llm_litellm_groq_llama_3.1_8b_instant.
Figure 12: Pie breakdown (tic_tac_toe): llm_litellm_groq_llama3_8b_8192.
Figure 13: Pie breakdown (kuhn_poker): llm_litellm_groq_llama3_70b_8192.
5.3.5 Entropy Across Agents Within a Game
To quantify cross-model diversity of reasoning over time for a fixed game, we compute the Shannon entropy of the reasoning-type distribution at each turn bin, aggregating across agents. Higher entropy indicates a larger spread across reasoning types, whereas zero entropy reflects unanimity.
Figure 14 reports entropy across agents for tic_tac_toe. The llm_litellm_groq_llama3_8b_8192 model shows a clear mid-game peak (greater strategic diversity), followed by a collapse to near-zero in late turns, consistent with convergence to Blocking. The other agents remain near zero throughout.
Figure 14: Entropy by turn across agents (tic_tac_toe). Higher values indicate more diverse reasoning distributions across agents at a given turn; zero denotes unanimity.
5.3.6 Synthesis
Taken together, these views reveal a coherent pattern. First, LLMs adapt their reasoning across games in ways aligned with game structure (radar, stacked bars). Second, they adapt within a single game: openings display planning/anticipation, mid-games emphasise board control, and endgames converge to defensive blocking (heatmaps and turn-stacked bars). Third, model size influences diversity—larger models tend to display richer, more balanced profiles, while smaller ones rely on a narrower set of rationales. Finally, entropy confirms that strategic diversity varies over time and across agents, peaking when the search space is most open and collapsing as terminal constraints dominate.
@@ 7 Related Research
Collectively, these benchmarks indicate that while LLMs excel at many language tasks, strategic gameplay remains challenging. They highlight the importance of dynamic, multi‑agent and interactive environments for evaluating reasoning, planning and social skills, and suggest motivation for frameworks such as Game Reasoning Arena
Appendix A Appendix
The architecture of our codebase is organized into modular components that separate concerns across environment design, agent control, and simulation execution. As illustrated in Figure 15, the simulation process begins with runner.py, which serves as the main entry point and initializes the overall configuration. This script delegates the core simulation logic to simulate.py, which orchestrates the interaction between the environment, game registry, and agent policies. On the left-hand side of the diagram, simulate.py interfaces with registry.py, a module responsible for loading and registering matrix games through the Matrix Game Loader. At the center of the architecture, environment classes such as MatrixGameEnv are defined to implement the game logic and extend a common OpenSpielEnv base class to ensure compatibility with the OpenSpiel interface. On the right-hand side, agent configuration and management are handled by policy_manager.py, which supports multiple agent types, including large language models (LLMs), random agents, and human-controlled agents. Each agent computes its actions based on the current game state and proceeds through the simulation loop via the action computation and game step execution phases. Finally, game outcomes and agent behaviors are recorded through a structured logging mechanism, enabling detailed analysis and post-hoc evaluation. This modular and extensible design facilitates efficient benchmarking of LLM strategies in matrix game settings.
Figure 15: High-level architecture of the Game Reasoning Arena framework for matrix games.
The interaction flow of a single matrix game episode within the Game Reasoning Arena is depicted in Figure 16. The simulation begins with environment initialization via env.reset(), followed by the construction of player-specific observations using the _state_to_observation() method. The simulation loop then enters the Action Computation Phase, where each agent selects an action based on its observation. For Player 0 (an LLM-controlled agent), this involves querying a language model using a prompt that includes the legal actions. The LLM returns both an action and its reasoning (e.g., “Paper beats Rock”). For Player 1 (a random agent), actions are sampled uniformly. In the Simultaneous Step Phase, the environment executes the joint action using env.step(), which internally calls state.apply_actions() to apply the moves and computes rewards via _compute_reward(). After each step, the environment returns updated observations, rewards, and a termination flag. Finally, during the Logging Phase, the system records all relevant information, including the LLM’s reasoning, action choices, and outcome metrics, using structured logs (e.g., SQLiteLogger) and real-time visualization tools such as TensorBoard. This loop continues until the game reaches a terminal state.
Figure 16: Sequence diagram of a single matrix game episode in the Game Reasoning Arena. It illustrates environment setup, LLM and random agent decision-making, game state updates, and structured logging.